# Explicit Context Reasoning with Supervision for Visual Tracking

### v1

@@ 1. Introduction
Visual object tracking (Bertinetto et al., 2016; Chen et al., 2022; Hu et al., 2025; Xue et al., 2024, 2025; Shi et al., 2025) is a fundamental task in computer vision. It aims to locate the object in subsequent frames using a bounding box, given the initial state of an arbitrary object in the first frame. However, in some complex tracking scenarios, such as fast motion and similar object interference or occlusion, traditional methods (Xu et al., 2020; Li et al., 2018; Xie et al., 2022) relying on static initial frame often struggle to handle dynamic target appearance variations or background interference, leading to poor robustness.
@@
To overcome the limitations of static references, researchers (Yan et al., 2021; Wei et al., 2023; Hu et al., 2024a; Zheng et al., 2024) have explored leveraging historical target features as priors to capture correlations with current frame, thereby enabling dynamic modeling of both the target and the scene. This topic has attracted widespread attention and developed rapidly, with mainstream methods transitioning from early discrete dynamic template updating (Yan et al., 2021; Cui et al., 2022; Fu et al., 2021) to current continuous video-level context modeling (Bai et al., 2024; Xie et al., 2024b; Li et al., 2024). Overall, current video-level context modeling methods can be broadly categorized into two types: 1) Contextual propagation: This approach (Zheng et al., 2024; Shi et al., 2024) passes lightweight features or tokens along the temporal dimension to establish temporal association between the initial and current frames. However, the methods that aggregate target information from multiple frames using unified features or tokens often overlook the distinctive features of the target in different frames, making it hard to capture the contextual association accurately and resulting in poor temporal modeling. 2) Contextual Association Reasoning: This method (Xie et al., 2024a, b) generates a unique target state token for each frame. The state tokens from multiple historical frames form a continuous sequence of target states, which is then modeled and inferred along the temporal dimension. However, due to the lack of effective supervision during this process, the model struggles to fully capture the temporal dynamics of the target state, resulting in modeling outputs that deviate significantly from the actual target state.
@@
To effectively learn and explicitly supervise the context reasoning process, we propose a novel context learning framework, RSTrack, aiming to model target consistency in video sequences to assist current state reason. As shown in Fig.1, we innovatively uses the true target state as a forward supervision signal to guide the continuous reasoning from historical to current states. Compared with existing implicit context modeling approaches (Zheng et al., 2024; Xie et al., 2024b; Li et al., 2024), RSTrack explicitly learns and leverages the temporal consistency among historical states to infer the current target state, thereby achieving context reasoning modeling with a clear evolutionary direction. Specifically, RSTrack achieves the above objectives through the following three components: 1) Efficient State Modeling: We propose a spatial-channel compression module that reduces redundant target features across frames by compressing them into state tokens that only retain essential object information. To ensure effective compression, we reconstruct the original features by combining the tokens with template features, and regularize the process using an L2 loss. This mechanism not only enables efficient information storage but also mitigates interference caused by inter-frame redundancy during context reasoning, thereby avoiding invalid contextual associations. 2) Context Reasoning Mechanism: We use a state space model to capture temporal variations of the target across frames, modeling the temporal correlations between historical target states to predict the current state. This predicted state token is converted to predicted target features via the reconstruction mechanism, which further refines the search feature in the temporal decoder. This mechanism fully leverages temporal consistency, enabling the model to reason more accurately about the current frame based on past states. 3) Forward Supervision Strategy: To constrain the context reasoning process, we compress the target features from the visual encoder into state tokens as supervision signals. By computing the L2 norm distance between the predicted state tokens and the supervision signals, we establish an explicit constraint mechanism that ensures effective learning of the contextual reasoning process and suppresses drift during the modeling process. In summary, we make the following contributions:
@@ 2.1. Temporal modeling for visual tracking
Traditional visual tracking methods (Bertinetto et al., 2016; Li et al., 2018; Hu et al., 2024b, c) usually rely on the initial template frame as a reference to track the target in subsequent search frames. This approach, which relies solely on the initial static appearance, often performs poorly when the target undergoes appearance variations or similar objects interference. To address the problem, researchers (Fu et al., 2021; Yan et al., 2021; Chen et al., 2023) have begun to introduce additional target references. For example, STARK (Yan et al., 2021) introduces dynamic templates and updates them using an additional score head; MixFormer (Cui et al., 2022) updates templates through fixed intervals and a score prediction module; STMTrack (Fu et al., 2021) stores historical target information in a memory network and uses the current frame as a query to retrieve target-related features from the memory bank. However, these discrete accumulation methods perform poorly when modeling the continuous variations of the target. Currently, many researchers (Zheng et al., 2024; Shi et al., 2024; Li et al., 2024) have focused on context propagation, using lightweight features or tokens to model cross-frame context. For example, ODTrack (Zheng et al., 2024) propagates historical tokens to establish cross-frame associations; EVPTrack (Shi et al., 2024) abstracts multi-scale features into explicit visual prompts to enhance contextual connections. However, these methods often overlook the uniqueness of features in different frames, as they aggregate historical information using unified features. Recent research, TemTrack (Xie et al., 2024a), generates unique target state tokens for each frame and models the sequence of historical frame state tokens along the temporal dimension. This approach preserves the uniqueness of the target in each frame, making the context modeling process more refined. However, without effective supervision, such implicit context learning may lead to model reasoning drift and fail to precisely capture the temporal cross-frame consistency of targets. In contrast, our approach introduces a forward supervision mechanism that explicitly models the dynamic changes in the target appearance across the video sequence, effectively suppressing drift in the reasoning process and ensuring better temporal consistency.
@@ 3. Method
In this section, we will provide a detailed explanation of the proposed RSTrack. First, we present a brief overview of the overall framework. Then, we describe each module of the whole framework. Finally, we introduce the training and inference pipelines.
@@ 3.1. Overview
The overall framework of RSTrack is shown in Fig.2. First, the template frame and search frame are embedded as 1D tokens and processed by a visual encoder to extract initial target features . Then, the state reasoning module models the historical state sequence to predict the current target state. Based on this prediction, the reconstruction module transforms the template features to generate predicted target features . Next, the temporal decoder refines the initial features using , producing the final temporally modeled features for the head network to locate the target. After each tracking iteration, the true target features are compressed and appended to to maintain the historical state sequence, while also serving as supervisory signals to guide reasoning and explicitly learn contextual changes.
@@ 3.3. Context Reasoning Mechanism
Here, denotes global pooling along the width or height dimension, represents depthwise separable 1D convolution with a kernel size of , indicates concatenating all along the channel dimension, represents Group Normalization with a group size of 4, and is the Sigmoid activation function. By accumulating compression results over multiple time steps, we can construct a history state sequence of the target. denoted as . The sequence is updated at each time step by appending the current state tokens to the end.
@@ Acknowledgements.
This work is supported by the Project of Guangxi Science and Technology (No.2025GXNSFAA069676, 2025GXNSFAA069417 and 2024GXNSFGA010001), the National Natural Science Foundation of China (No.U23A20383, 62472109, 62466051 and 62402252), the Guangxi “Young Bagui Scholar” Teams for Innovation and Research Project, the Research Project of Guangxi Normal University (No.2024DF001), the grant from Guangxi Colleges and Universities Key Laboratory of Intelligent Software (No.2024B01).References

### v2

@@ 1. Introduction
Visual object tracking (Bertinetto et al., 2016; Chen et al., 2022; Hu et al., 2025; Xue et al., 2024, 2025; Shi et al., 2025) is a fundamental task in computer vision. It aims to locate the object in subsequent frames using a bounding box, given the initial state of an arbitrary object in the first frame. However, in some complex tracking scenarios, such as fast motion and similar object interference or occlusion, traditional methods (Xu et al., 2020; Li et al., 2018; Xie et al., 2022) relying on a static initial frame often struggle to handle dynamic target appearance variations or background interference, leading to poor robustness.
@@
To overcome the limitations of static references, researchers (Yan et al., 2021; Wei et al., 2023; Hu et al., 2024a; Zheng et al., 2024) have explored leveraging historical target features as priors to capture correlations with the current frame, thereby enabling dynamic modeling of both the target and the scene. This topic has attracted widespread attention and developed rapidly, with mainstream methods transitioning from early discrete dynamic template updating (Yan et al., 2021; Cui et al., 2022; Fu et al., 2021) to current continuous video-level context modeling (Bai et al., 2024; Xie et al., 2024b; Li et al., 2024). Overall, current video-level context modeling methods can be broadly categorized into two types: 1) Contextual Propagation: This approach (Zheng et al., 2024; Shi et al., 2024) passes lightweight features or tokens along the temporal dimension to establish temporal association between the initial and current frames. However, the methods that aggregate target information from multiple frames using unified features or tokens often overlook the distinctive features of the target in different frames, making it hard to capture the contextual association accurately and resulting in poor temporal modeling. 2) Contextual Association Reasoning: This method (Xie et al., 2024a, b) generates a unique target state token for each frame. The state tokens from multiple historical frames form a continuous sequence of target states, which is then modeled and inferred along the temporal dimension. However, due to the lack of effective supervision during this process, the model struggles to fully capture the temporal dynamics of the target state, resulting in modeling outputs that deviate significantly from the actual target state.
@@
To effectively learn and explicitly supervise the context reasoning process, we propose a novel context learning framework, RSTrack, aiming to model target consistency in video sequences to assist current state reasoning. As shown in Fig.1, we innovatively use the true target state as a forward supervision signal to guide the continuous reasoning from historical to current states. Compared with existing implicit context modeling approaches (Zheng et al., 2024; Xie et al., 2024b; Li et al., 2024), RSTrack explicitly learns and leverages the temporal consistency among historical states to infer the current target state, thereby achieving context reasoning modeling with a clear evolutionary direction. Specifically, RSTrack achieves the above objectives through the following three components: 1) Efficient State Modeling: We propose a spatial-channel compression module that reduces redundant target features across frames by compressing them into state tokens that only retain essential object information. To ensure effective compression, we reconstruct the original features by combining the tokens with template features, and regularize the process using an L2 loss. This mechanism not only enables efficient information storage but also mitigates interference caused by inter-frame redundancy during context reasoning, thereby avoiding invalid contextual associations. 2) Context Reasoning Mechanism: We use a state space model to capture temporal variations of the target across frames, modeling the temporal correlations between historical target states to predict the current state. This predicted state token is converted to predicted target features via the reconstruction mechanism, which further refines the search feature in the temporal decoder. This mechanism fully leverages temporal consistency, enabling the model to reason more accurately about the current frame based on past states. 3) Forward Supervision Strategy: To effectively constrain the context reasoning process, we compress the target features extracted from the visual encoder into compact state tokens, which serve as reliable supervision signals. By computing the L2 norm distance between the predicted state tokens and the supervision signals, we establish an explicit constraint mechanism that ensures effective learning of the contextual reasoning process and suppresses drift during the modeling process. In summary, we make the following contributions:
@@ 2.1. Temporal modeling for visual tracking
Traditional visual tracking methods (Bertinetto et al., 2016; Tang et al., 2024; Hu et al., 2024b, c; Tang et al., 2023) usually rely on the initial template frame as a reference to track the target in subsequent search frames. This approach, which depends solely on the initial static appearance, often performs poorly when the target undergoes significant appearance variations or similar object interference. To address the problem, researchers (Fu et al., 2021; Yan et al., 2021; Chen et al., 2023) have increasingly begun to introduce additional target references. For example, STARK (Yan et al., 2021) introduces dynamic templates and updates them using an additional score head; MixFormer (Cui et al., 2022) updates templates through fixed intervals and a score prediction module; STMTrack (Fu et al., 2021) stores historical target information within a memory network and utilizes the current frame as a query to retrieve highly relevant target features from the memory bank. However, these discrete accumulation methods generally perform poorly when modeling the continuous variations of the target. More recently, many researchers (Zheng et al., 2024; Shi et al., 2024; Li et al., 2024) have focused on context propagation, using lightweight features or tokens to model cross-frame context. For example, ODTrack (Zheng et al., 2024) propagates historical tokens to establish cross-frame associations; EVPTrack (Shi et al., 2024) abstracts multi-scale features into explicit visual prompts to enhance contextual connections. Nevertheless, these methods often overlook the uniqueness of features in different frames, as they aggregate historical information using unified features. Recent research, TemTrack (Xie et al., 2024a), generates unique target state tokens for each frame and models the sequence of historical frame state tokens along the temporal dimension. This approach preserves the uniqueness of the target in each frame, making the context modeling process more refined. However, without effective supervision, such implicit context learning may lead to model reasoning drift and fail to precisely capture the temporal cross-frame consistency of targets. In contrast, our approach introduces a forward supervision mechanism that explicitly models the dynamic changes in the target appearance across the video sequence, effectively suppressing drift in the reasoning process and ensuring better temporal consistency.
@@ 3. Method
In this section, we will provide a detailed explanation of the proposed RSTrack. First, we present a brief overview of the overall framework. Then, we describe each module in detail to explain its roles and interactions. Finally, we introduce the training and inference pipelines to outline the complete workflow.
@@ 3.1. Overview
The overall framework of RSTrack is shown in Fig.2. First, the template frame and search frame are embedded as 1D tokens and processed by a visual encoder to extract initial target features . Subsequently, the state reasoning module uses the historical state sequence to predict the current target state. Based on this prediction, the reconstruction module transforms the template features to generate predicted target features . Next, the temporal decoder refines the initial search features using , producing the final temporally enriched features for the head network to locate the target. After each iteration, the true target features are compressed and appended to , maintaining the continuity of the historical states. Additionally, these true features , along with the compressed state tokens, serve as supervisory signals to guide reasoning and explicitly capture contextual dynamics.
@@ 3.3. Context Reasoning Mechanism
Here, denotes global pooling along the width or height dimension, represents depthwise separable 1D convolution with a kernel size of , indicates concatenating all along the channel dimension, represents Group Normalization with a group size of 4, and is the Sigmoid activation function. By accumulating compression results over multiple time steps, we can construct a history state sequence of the target. denoted as At each time step, is updated by appending the current state tokens to the end, maintaining the target’s temporal states.
@@ Acknowledgements.
This work is supported by the Project of Guangxi Science and Technology (No.2025GXNSFAA069676 and 2024GXNSFGA010001), the National Natural Science Foundation of China (No.U23A20383, 62472109, 62466051 and 62402252), the Guangxi “Young Bagui Scholar” Teams for Innovation and Research Project, the Research Project of Guangxi Normal University (No.2024DF001), the grant from Guangxi Colleges and Universities Key Laboratory of Intelligent Software (No.2024B01).References
@@
- Tang et al. (2023) Zhangyong Tang, Tianyang Xu, Hui Li, Xiao-Jun Wu, Xuefeng Zhu, and Josef Kittler. 2023. Exploring fusion strategies for accurate RGBT visual object tracking. Information Fusion 99 (2023), 101881.
- Tang et al. (2024) Zhangyong Tang, Tianyang Xu, Xiaojun Wu, Xue-Feng Zhu, and Josef Kittler. 2024. Generative-based fusion mechanism for multi-modal tracking. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 5189–5197.